A line with slope 1/5 passes through (6,3). If the abscissa of a point on the line is 11, what is the ordinate? NEED HELP. THANKS IN ADVANCE
Answer by richwmiller(17219)

y-3=1/5*(x-6)
y = x/5+9/5